# Creating and Benchmarking a Synthetic Dataset for Cloud Optical Thickness Estimation

### v1

@@ Abstract
Cloud formations often obscure optical satellite-based monitoring of the Earth’s surface, thus limiting Earth observation (EO) activities such as land cover mapping, ocean color analysis, and cropland monitoring. The integration of machine learning (ML) methods within the remote sensing domain has significantly improved performance on a wide range of EO tasks, including cloud detection and filtering, but there is still much room for improvement. A key bottleneck is that ML methods typically depend on large amounts of annotated data for training, which is often difficult to come by in EO contexts. This is especially true for the task of cloud optical thickness (COT) estimation. A reliable estimation of COT enables more fine-grained and application-dependent control compared to using pre-specified cloud categories, as is commonly done in practice. To alleviate the COT data scarcity problem, in this work we propose a novel synthetic dataset for COT estimation, where top-of-atmosphere radiances have been simulated for 12 of the spectral bands of the Multi-Spectral Instrument (MSI) sensor onboard Sentinel-2 platforms. These data points have been simulated under consideration of different cloud types, COTs, and ground surface and atmospheric profiles. Extensive experimentation of training several ML models to predict COT from the measured reflectivity of the spectral bands demonstrates the usefulness of our proposed dataset. Generalization to real data is also demonstrated on two satellite image datasets – one that is publicly available, and one which we have collected and annotated. The synthetic data, the newly collected real dataset, code and models have been made publicly available at https://github.com/aleksispi/ml-cloud-opt-thick.
@@ 1 Introduction
The cloud detection and estimation solutions in multi-spectral images range from rule-based statistical methods to advanced deep learning approaches. On the one hand, rule-based thresholding methods exploit the physical properties of clouds reflected on different spectral bands of satellite images to generate cloud masks. FMask [33] and Sen2Cor [20] are examples of such approaches implemented for cloud masking for Landsat and Sentinel-2 multi-spectral imagery, respectively. On the other hand, recently a burst of work in the literature has been using ML approaches to solve the cloud detection and estimation problem [22, 32, 15, 14]. These works use convolutional neural networks with large manually annotated datasets of satellite images and perform better than statistical rule-based methods.
Clouds are inhomogeneous by nature and the spatial inhomogeneity of clouds, or their varying cloud optical thickness (COT), affects not only the remote sensing imagery but also atmospheric radiations. Most literature estimates COT using independent pixel analysis (IPA, also known as ICA, independent column approximation [27]). IPA considers the cloud properties homogeneous in the pixel and carries no information about the neighboring pixels. Statistical approaches  [19, 34, 13] have explored the potential causes influencing COT and defined parameters to mitigate their effect. With the advancement of deep learning [18, 11], some approaches for COT using neural networks have been proposed [24, 28]. Most of these methods require neighboring information of the pixels and large amounts of annotated spatial data.
@@
In this work we propose a new IPA dataset simulating top-of-atmosphere (TOA) radiances for 12 of the spectral bands [^1] The aerosol (B1) band is not included in the simulation, as it adds another dimension of computational complexity to the simulation (its inclusion is an avenue for future work) of the Multi Spectral Imager (MSI) onboard the Sentinel-2A platform. We are aware that the IPA treatment of the data, especially for high spacial resolution applications, will introduce a systematic error in COT, due to 3D effects [35]. Nevertheless, the derivation of COT allows for more flexible use of data (e.g. clear-conservative vs cloud-conservative cloud-mask). The dataset has been generated using RTTOV v13 [26] and is composed of 200,000 data points. We present results on a range of ML models trained for cloud detection and COT estimation using our dataset. The experimental results suggest that on unseen data points, multi-layer perceptrons (MLPs) significantly outperform alternative approaches, like linear regression. We have also investigated the performance of our trained models on two real satellite image datasets – the publicly available KappaZeta, [7] as well as a novel dataset provided by the Swedish Forest Agency. These results show that models trained on our proposed synthetic dataset generalize to real data and use-cases.
@@ 2 Synthetic Cloud Optical Thickness Dataset
In this section we explain the synthetic dataset that we propose for the task of COT estimation. The dataset consists of 200,000 simulated data points (akin to individual and independent pixels, as observed by a satellite instrument), which have been simulated taking into consideration different cloud types, cloud optical thicknesses (COTs), cloud geometrical thickness, cloud heights, as well as ground surface and atmospheric profiles. Variations in water vapour content, gaseous optical thickness, and angles (azimuth difference as well as satellite and solar zenith angle) have also been included in the dataset.
@@
Figure 1: Distribution of COTs in the range $[0,50]$ for different cloud types in our dataset: water clouds (left), ice clouds (center), and combined water and ice clouds (right). The focus of our data lies on optically thin clouds. Note that the frequency bins will not sum up to 50,000 (see main text for details).
The backbone of each simulation is an atmospheric profile, defined by the vertical distribution of pressure, temperature, water vapor content and ozone. These atmospheric profiles are picked at random from an ECMWF (European Centre for Medium-range Weather Forecasts) dataset of 10,000 profiles compiled with the aim of mapping the widest possible range of atmospheric conditions [6]. The surface is assumed to work as a lambertian reflector [17]. All spectral reflectance are provided by the ECOSTRESS spectral library [3, 23].
Each main surface type is composed of several sub-groups. The main surface types are non-photosynthetic vegetation (e.g. tree bark, branches, flowers), rock (e.g. gneis, marble, limestone), vegetation (e.g. grass, shrub, tree), soil (e.g. calciorthid, quartzipsamment, cryoboroll) and water (ice, snow in different granularities, and liquid). The total number of different reflectivity profiles (including all not listed sub-groups) comprises 743 variations. These 743 fine classes appear in the dataset as 139 distinguishable categories. When generating the data, surface types were sampled at random, and the distribution of the main types within the final dataset is shown in Table 1. The instrument-specific reflection was derived by convolving the spectral surface reflection with the spectral response function of the corresponding MSI channel.
The dataset consists of four equal-sized parts (50,000 points each). One part features clear sky calculations only (no clouds are considered here, only atmospheric gases and surface properties), while the three remaining parts respectively depict water clouds, ice clouds, and mixed clouds. Water clouds are parameterized following the size distributions of the five OPAC cloud types [12] with updates in the liquid water refractive index dataset according to [29]. Ice clouds are parameterized by temperature and ice water content (IWC) according to [4]. Distributions of COT for water, ice and combined clouds are shown in Fig. 1, where it can be seen that our focus is set on optically thin clouds. Note that when a consistency-check fails, the cloud will be removed and the COT is set to 0, but the data point remains in its place. Thus, within the dataset there are some data points that are labeled cloudy even though they depict clear sky cases. These clear sky cases are removed from the plots in Fig. 1.
@@
where we successfully perform classification of real image pixels into semi-transparent and opaque (optically thin and thick) clouds. We do this even though we understand that the pure prediction of the optical density, due to 3D effects, has little significance. Note that according to the international satellite cloud climatology project (ISCCP, see e.g. [25]), clouds are considered optically thin for $\text{COT}<3.6$, medium for $3.6\leq\text{COT}<23$ and optically thick for $\text{COT}\geq 23$. Keeping in mind that optically thick clouds are in general easier to detect, it is not expected that this limitation affects our conclusions.
To summarize, the proposed dataset $\mathcal{D}=\{(\boldsymbol{x}_{1},y_{1}),\dots,\boldsymbol{x}_{N},y_{n})\}$ consists of 200,000 pairs $(\boldsymbol{x}_{i},y_{i})$ with $\boldsymbol{x}_{i}\in\mathbb{R}^{19}$ and $y_{i}\in\mathbb{R}_{+}$ the $i$:th input and ground truth COT (non-negative), respectively. Here, $\boldsymbol{x}_{i}$ contains simulated measured reflectivities for the 12 spectral bands, but also additional optional features: (i) satellite zenith angle [${}^{\circ}$]; (ii) sun zenith angle [${}^{\circ}$]; (iii) azimuth difference angle [${}^{\circ}$]; (iv) gas optical thickness [--]; (v) vertically integrated water vapour [$\text{g}/\text{cm}^{2}$]; (vi) surface profile [--]; and (vii) cloud type [--]. In this work we have focused on experiments using the 12 band reflectivities as model input, which are least effected by aerosols, and will thus override notation and let $\boldsymbol{x}_{i}\in\mathbb{R}^{12}$ denote such a data point. As for the ground truth COT $y_{i}$, these lie in the range $y_{i}\in[0,50]$; again, the upper limit is set by the radiative transfer model. We randomly set aside 160,000 data points for model training, 20,000 for validation, and 20,000 for testing (see
@@ 8 Machine Learning Models
In this section we briefly introduce the different machine learning (ML) models that we have implemented for the COT estimation task. As our proposed dataset can be seen as consisting of individual pixels, with no spatial relationship in-between them (cf.
@@
), we have mainly considered MLP models. [^2] Early experiments were done with a range of approaches, e.g. Random Forests, but MLPs performed best. Extensive model validations suggested that a five-layer MLP with $\mathrm{ReLU}{}$-activations and hidden dimension $64$ (same for all layers) yields the best results.
When performing inference on real imagery $\boldsymbol{I}\in\mathbb{R}^{H\times W\times C}$ with an MLP model, we let it operate at a per-pixel basis, as the models are trained (see
@@
) on individual synthetic pixels due to the nature of the synthetic data. To improve spatial consistency of a resulting COT prediction map $\boldsymbol{C}\in\mathbb{R}^{H\times W}$, we apply a simple post-processing scheme where a sliding window of size $M\times M$ and stride 1 runs on top of $\boldsymbol{C}\in\mathbb{R}^{H\times W}$, and produces an average (among $M^{2}$ values) at each location. Overlaps that occur due to the stride being smaller than the width of the kernel are handled by averaging. We found $M=2$ to produce robust results.
@@ 10.1 Model Training
We apply standardization (to channel-wise zero mean and unit variance, based on training data statistics) to the input, as more sophisticated data normalization techniques did not yield improvements. The ML models are trained using the Adam optimizer [16] with an L2-loss for 2e6 batch updates (batch size 32, learning rate 3e-4).
To improve model robustness, independent Gaussian noise is added to each input. This noise is zero-mean and has a standard deviation corresponding to 3% of the average magnitude of each input feature. Since the models are very lightweight, training can be done even without a GPU within about an hour.

### v3

# Creating and Leveraging a Synthetic Dataset of Cloud Optical Thickness Measures for Cloud Detection in MSI

@@ Abstract
Cloud formations often obscure optical satellite-based monitoring of the Earth’s surface, thus limiting Earth observation (EO) activities such as land cover mapping, ocean color analysis, and cropland monitoring. The integration of machine learning (ML) methods within the remote sensing domain has significantly improved performance on a wide range of EO tasks, including cloud detection and filtering, but there is still much room for improvement. A key bottleneck is that ML methods typically depend on large amounts of annotated data for training, which is often difficult to come by in EO contexts. This is especially true when it comes to cloud optical thickness (COT) estimation. A reliable estimation of COT enables more fine-grained and application-dependent control compared to using pre-specified cloud categories, as is commonly done in practice. To alleviate the COT data scarcity problem, in this work we propose a novel synthetic dataset for COT estimation, that we subsequently leverage for obtaining reliable and versatile cloud masks on real data. In our dataset, top-of-atmosphere radiances have been simulated for 12 of the spectral bands of the Multispectral Imagery (MSI) sensor onboard Sentinel-2 platforms. These data points have been simulated under consideration of different cloud types, COTs, and ground surface and atmospheric profiles. Extensive experimentation of training several ML models to predict COT from the measured reflectivity of the spectral bands demonstrates the usefulness of our proposed dataset. In particular, by thresholding COT estimates from our ML models, we show on two satellite image datasets (one that is publicly available, and one which we have collected and annotated) that reliable cloud masks can be obtained. The synthetic data, the newly collected real dataset, code and models have been made publicly available at https://github.com/aleksispi/ml-cloud-opt-thick.
@@ 1 Introduction
The cloud detection and estimation solutions in multispectral images range from rule-based statistical methods to advanced deep learning approaches. On the one hand, rule-based thresholding methods exploit the physical properties of clouds reflected on different spectral bands of satellite images to generate cloud masks. FMask [33] and Sen2Cor [20] are examples of such approaches implemented for cloud masking for Landsat and Sentinel-2 multispectral imagery, respectively. On the other hand, recently a burst of work in the literature has been using ML approaches to solve the cloud detection and estimation problem [22, 32, 15, 14]. These works use convolutional neural networks with large manually annotated datasets of satellite images and perform better than statistical rule-based methods.
Clouds are inhomogeneous by nature and the spatial inhomogeneity of clouds, or their varying cloud optical thickness (COT), affects not only the remote sensing imagery but also atmospheric radiation. The COT is an excellent proxy for a versatile cloud mask. By choosing a threshold value for the cloud/clear distinction (or even finer-grained categories), both a cloud-conservative and a clear-conservative cloud mask can be implemented. Most literature estimates COT using independent pixel analysis (IPA, also known as ICA, independent column approximation [27]). IPA considers the cloud properties homogeneous in the pixel and carries no information about the neighboring pixels. Statistical approaches  [19, 34, 13] have explored the potential causes influencing COT and defined parameters to mitigate their effect. With the advancement of deep learning [18, 11], some approaches for COT using neural networks have been proposed [24, 28]. Most of these methods require neighboring information of the pixels and large amounts of annotated spatial data.
@@
1.1 Contributions
In the below, we provide details of the contributions of this work, which can be summarized as (i) the creation and public release of a synthetic 1D dataset simulating top-of-atmosphere (TOA) radiances for 12 of the spectral bands of the Multispectral Imagery (MSI) onboard the Sentinel-2A platform; and (ii) an extensive set of experiments with different ML approaches, which show that COT estimation can be used as a flexible proxy for obtaining reliable and versatile cloud masks on real data.
A new synthetic 1D dataset of TOA radiances for MSI. We propose a new 1D dataset simulating TOA radiances for 12 of the spectral bands [^1] The aerosol (B1) band is not included in the simulation, as it adds another dimension of computational complexity to the simulation (its inclusion is an avenue for future work) of the MSI onboard the Sentinel-2A platform. The data points have been simulated taking into consideration different cloud types, cloud optical thicknesses (COTs), cloud geometrical thickness, cloud heights, water vapour contents, gaseous optical thicknesses, as well as ground surface and atmospheric profiles. For this, we connected the resources natively available in RTTOV v13 [26] with external resources, such as a dataset of atmospheric profiles provided by ECMWF, and a dataset of spectral reflectances from the ECOSTRESS spectral library [3, 23]. More details about the dataset are provided in
2 sec:synth-data
The public release of this data implies that the threshold of research participation is lowered, in particular for non-domain experts who wish to contribute to this field, and offers reproducible and controllable benchmarking of various methods and approaches. Potential users of our dataset should be aware of the possible limitations the IPA could introduce in their study. In that regard, we note that if the users are interested in developing high spacial resolution applications, IPA will introduce a systematic error in COT, due to 3D effects [35]. Nevertheless, the derivation of COT allows for more flexible use of data (e.g. clear-conservative vs cloud-conservative cloud-mask).
Extensive ML model experimentation based on synthetic and real data. We present results on a range of ML models trained for cloud detection and COT estimation using our proposed dataset. The experimental results suggest that on unseen data points, multi-layer perceptrons (MLPs) significantly outperform alternative approaches, like linear regression. We have also investigated and demonstrated the performance of our trained models on two real satellite image datasets – the publicly available KappaZeta, [7] as well as a novel dataset provided by the Swedish Forest Agency (which has also been made publicly available). These results show that by thresholding COT estimates from our ML models trained on the proposed synthetic dataset, reliable cloud masks can be obtained on real images; they even outperform the scene classification layer by ESA on the dataset provided by the Swedish Forest Agency. The main ML models are described in
3 sec:models
and the experimental results are included in
4 sec:results
@@ 5 Synthetic Cloud Optical Thickness Dataset
In this section we explain the COT synthetic dataset that we propose to use for subsequent threshold-based cloud masking. The dataset consists of 200,000 simulated data points (akin to individual and independent pixels, as observed by a satellite instrument), which have been simulated taking into consideration different cloud types, cloud optical thicknesses (COTs), cloud geometrical thickness, cloud heights, as well as ground surface and atmospheric profiles. Variations in water vapour content, gaseous optical thickness, and angles (azimuth difference as well as satellite and solar zenith angle) have also been included in the dataset.
@@
The backbone of each simulation is an atmospheric profile, defined by the vertical distribution of pressure, temperature, water vapor content and ozone. These atmospheric profiles are picked at random from an ECMWF (European Centre for Medium-range Weather Forecasts) dataset of 10,000 profiles compiled with the aim of mapping the widest possible range of atmospheric conditions [6]. The surface is assumed to work as a lambertian reflector [17]. All spectral reflectance are provided by the ECOSTRESS spectral library [3, 23]. Sun and satellite angles are varied randomly.
@@
Each main surface type is composed of several sub-groups. The main surface types are non-photosynthetic vegetation (e.g. tree bark, branches, flowers), rock (e.g. gneis, marble, limestone), vegetation (e.g. grass, shrub, tree), soil (e.g. calciorthid, quartzipsamment, cryoboroll) and water (ice, snow in different granularities, and liquid). The total number of different reflectivity profiles (including all not listed sub-groups) comprises 743 variations. These 743 fine classes appear in the dataset as 139 distinguishable categories. The distribution of the main types within the final dataset is shown in Table 1. The instrument-specific reflection was derived by convolving the spectral surface reflection with the spectral response function of the corresponding MSI channel. When generating the data, surface types were sampled at random.
Figure 1: Distribution of COTs in the range $[0,50]$ for different cloud types in our dataset: water clouds (left), ice clouds (center), and combined water and ice clouds (right). The focus of our data is on optically thin clouds.
The dataset consists of four equal-sized parts (50,000 points each). One part features clear sky calculations only (no clouds are considered here, only atmospheric gases and surface properties), while the three remaining parts respectively depict water clouds, ice clouds, and mixed clouds. Cloud geometrical thickness is set by random. The clouds geometrical thickness varies from 1 layer to 6 with the mode at 2 and 3. Clouds of the same type populate adjacent layers but vertical gaps are possible between ice and water clouds. The cloud type (for water clouds) and liquid water/ice content is varied randomly. Water clouds are parameterized following the size distributions of the five OPAC cloud types [12] with updates in the liquid water refractive index dataset according to [29]. Ice clouds are parameterized by temperature and ice water content (IWC) according to [4]. Variations in size and shape distributions are taken into account implicitly by the parametrization of particle optical properties. Distributions of COT for water, ice and combined clouds are shown in Fig. 1, where it can be seen that our focus is set on optically thin clouds. Note that when a consistency-check fails (i.e. the cloud type and background atmospheric temperature do not match), the cloud will be removed and the COT is set to 0, but the data point remains in its place. Thus, within the dataset there are some data points that are labeled cloudy even though they depict clear sky cases. These clear sky cases are removed from the plots in Fig. 1.
@@
where we successfully perform classification of real image pixels into semi-transparent and opaque (optically thin and thick) clouds. We do this even though we understand that the pure prediction of the optical density, due to 3D effects, has little significance. The most relevant 3D effect for MSI with a high spatial resolution is the neglect of the net horizontal radiation transport. The error depends on the horizontal component of the photon movement and thus largely on the solar zenith angle. However, even the incorrect COT is suitable for a cloudy/cloud-free decision. Note that according to the international satellite cloud climatology project (ISCCP, see e.g. [25]), clouds are considered optically thin for $\text{COT}<3.6$, medium for $3.6\leq\text{COT}<23$ and optically thick for $\text{COT}\geq 23$. Keeping in mind that optically thick clouds are in general easier to detect, it is not expected that this limitation affects our conclusions.
Figure 2: Main MLP model architecture that we use for COT estimation. The input layer is shown on the left (4 nodes shown for the input $\boldsymbol{x}$ instead of all 12, to avoid visual clutter). Then follows four hidden layers (8 nodes shown for each, instead of 64). Finally, the output layer on the right produces the COT estimate $\hat{y}$. Each of the five layers is of the form $\boldsymbol{z}^{k}=\mathrm{ReLU}{}(\boldsymbol{W}^{k}\boldsymbol{z}^{k-1}+\boldsymbol{b}^{k})=\mathrm{max}(\boldsymbol{0},\boldsymbol{W}^{k}\boldsymbol{z}^{k-1}+\boldsymbol{b}^{k})$ for $k=1,\dots,5$, with $\boldsymbol{z}^{0}=\boldsymbol{x}$ and $\boldsymbol{z}^{5}=\hat{y}$. Here, $\boldsymbol{W}^{k}$ and $\boldsymbol{b}^{k}$ respectively represent the weight matrix and bias vector for the $k$:th layer, and these parameters are calibrated through the training process (see
10 sec:training
). Note that we use an MLP, rather than e.g. a convolutional architecture, since the data points $\boldsymbol{x}_{i}$ are independent from each other in our synthetic dataset.
In summary, our dataset $\mathcal{D}=\{(\boldsymbol{x}_{1},y_{1}),\dots,\boldsymbol{x}_{n},y_{n})\}$ consists of 200,000 pairs $(\boldsymbol{x}_{i},y_{i})$ with $\boldsymbol{x}_{i}\in\mathbb{R}^{19}$ and $y_{i}\in\mathbb{R}_{+}$ the $i$:th input and ground truth COT, respectively, i.e. each element in the 19-dimensional vector $\boldsymbol{x}_{i}$ is a real scalar, and $y_{i}$ is a non-negative real scalar (since COT cannot be negative). By ground truth COT $y_{i}$, we mean the true COT against which COT predictions can be compared (e.g. measure how accurately a machine learning model predicts $y_{i}$ if it gets $\boldsymbol{x}_{i}$ as input). As for $\boldsymbol{x}_{i}$, it contains simulated measured reflectivities for the 12 spectral bands, but also additional optional features: (i) satellite zenith angle [${}^{\circ}$]; (ii) sun zenith angle [${}^{\circ}$]; (iii) azimuth difference angle [${}^{\circ}$]; (iv) gas optical thickness [--]; (v) vertically integrated water vapour [$\text{g}/\text{cm}^{2}$]; (vi) surface profile [--]; and (vii) cloud type [--].
In this work we have focused on experiments using only the 12 band reflectivities as model input, which are least effected by aerosols, and thus from this point onwards we will let $\boldsymbol{x}_{i}\in\mathbb{R}^{12}$ denote such a data point. One of the main reasons for not using other types of input information besides the band reflectivities is compatibility (to ensure that the resulting models would be able to be evaluated on publicly available real image datasets, see e.g.
11 sec:kappazeta
). As for the ground truth COT $y_{i}$, these lie in the range $y_{i}\in[0,50]$; again, the upper limit is set by the radiative transfer model. We randomly set aside 160,000 data points for model training, 20,000 for validation, and 20,000 for testing (see
@@ 14 Machine Learning Models
In this section we briefly introduce the different machine learning (ML) models that we have implemented for the COT estimation task. Recall from
@@
that the model inputs $\boldsymbol{x}_{i}\in\mathbb{R}^{12}$ contain only 12 band reflectivities, i.e. no auxiliary inputs such as surface profile information is used by the models. As our proposed dataset can be seen as consisting of individual ’pixels’, with no spatial relationship in-between them (cf.
16 sec:synth-data
), we have mainly considered multi-layer perceptron (MLP) models. [^2] Early experiments were done with a range of approaches, e.g. Random Forests, but MLPs performed best. Extensive model validations suggested that a five-layer MLP with $\mathrm{ReLU}{}$-activations and hidden dimension $64$ (same for all layers) yields the best results. See Fig. 10 for an overview of this architecture.
When performing inference on real imagery $\boldsymbol{I}\in\mathbb{R}^{H\times W\times C}$ with an MLP model, we let it operate at a pixel-per-pixel basis, as the models are trained (see
@@
) on individual synthetic pixels due to the nature of the synthetic data. To improve spatial consistency of a resulting COT prediction map $\boldsymbol{C}\in\mathbb{R}^{H\times W}$, we apply a simple post-processing scheme where a sliding window of size $M\times M$ and stride 1 runs on top of $\boldsymbol{C}\in\mathbb{R}^{H\times W}$, and produces an average (among $M^{2}$ values) at each location. Overlaps that occur due to the stride being smaller than the width of the kernel are handled by averaging. We found $M=2$ to produce robust results. Thus, via this post-processing each value within $\boldsymbol{C}$ is essentially replaced by the mean value of the surrounding values, which increases the smoothness of the predictions (nearby predictions are less likely to differ drastically). This is desirable, because in real images adjacent pixels are typically quite similar to each other, so the predictions for adjacent pixels are expected to be quite similar as well.
@@ 17.1 Model Training
As is standard within machine learning, we normalize the model inputs $\boldsymbol{x}_{i}\in\mathbb{R}^{12}$ such that each of the 12 entries of $\boldsymbol{x}_{i}$ has zero mean and unit standard deviation. This is done by first computing the mean $\boldsymbol{x}^{\text{mean}}\in\mathbb{R}^{12}$ and standard deviation $\boldsymbol{x}^{\text{std}}\in\mathbb{R}^{12}$ on the training set, and then for each $\boldsymbol{x}_{i}$ applying the following transformation (the subtraction and division below is done element-wise for each of the 12 entries) before feeding the result to the model:
$\boldsymbol{x}_{i}\leftarrow\frac{\boldsymbol{x}_{i}-\boldsymbol{x}^{\text{mean}}}{\boldsymbol{x}^{\text{std}}}.$ (1)
We also explored other data normalization techniques but did not see any improvements relative to (1). To improve model robustness, independent Gaussian noise is also added to each input. This noise is zero-mean and has a standard deviation corresponding to 3% of the average magnitude of each input feature. Thus, the transformation (1) is replaced by
$\boldsymbol{x}_{i}\leftarrow\frac{(\boldsymbol{x}_{i}+\boldsymbol{\varepsilon}_{i})-\boldsymbol{x}^{\text{mean}}}{\boldsymbol{x}^{\text{std}}},$ (2)
where $\boldsymbol{\varepsilon}_{i}\in\mathbb{R}^{12}$ denotes such an independently sampled Gaussian.
The ML models are trained with the commonly used Adam optimizer [16] with a mean squared error (MSE) loss for 2,000,000 batch updates (batch size 32, learning rate 0.0003). Since the models are very lightweight, training can be done even without a graphics processing unit (GPU) within about an hour.